# A Spectral Localization Method for Time-Fractional Integro-Differential Equations with Nonsmooth Data

### v1

@@ 4.1 Scalar Problem
$\begin{cases}{}_{0}^{C}D_{t}^{\beta}u(t)+u(t)+\int_{0}^{t}\kappa_{\alpha}(t-s)u(s)\,ds=f(t),\quad t>0,\\ u(0)=1,\end{cases}$ (4.1)
@@
(a) The absolute errors for different $\Lambda$ at $t=0.5$ with $\alpha=0.2$, $\beta=0.77$
(b) The absolute errors at different $t$ for $\Lambda=10$ with $\alpha=0.2$, $\beta=0.77$
(c) The absolute errors for different $\alpha$ at $t=0.5$ with $\beta=0.77$, $\Lambda=10$
(d) The absolute errors for different $\beta$ at $t=0.5$ with $\alpha=0.2$, $\Lambda=10$
Figure 3: Numerical performances of the CIM for problem (4.1)
@@ 4.2 Homogeneous Problems with Nonsmooth Initial Values in 1-D and 2-D
Consider the homogeneous problem with nonsmooth initial value for problem (1.1). In the 1-D case, the initial value is given by
@@
with source term $f(x,t)=0$. In the 2-D case, the initial value is taken as
@@
and the source term is $f(x,y,t)=0$.
To evaluate the numerical performance, we use the numerical solution computed with $N=200$ and $h=1/2^{7}$ as the reference solution. The numerical results at $t=0.4$ are presented in Tables 1-5 and Figures 5. Additionally, Figures 5 and 6 illustrate the behaviors of reference solution at different times, which also confirm that the evolutionary behavior of the solution is influenced by the regularity of the initial condition, as discussed in the Solution Theory section. For a better observation of the solution’s variation, in these two figures, the parameters are taken as $t_{0}=0.001$ and $\Lambda=1000$.
Table 1: Temporal errors $\mathrm{Err}_{\tau}(N)$ at $t=0.4$ for the 1-D homogeneous problem with nonsmooth initial data (4.2), shown for different values of $\alpha$ and $\beta$ with $h=1/2^{7}$ and $\Lambda=10$
@@
Table 2: Spatial errors $\mathrm{Err}_{h}$ at $t=0.4$ for the 1-D homogeneous problem with nonsmooth initial data (4.2), computed for different values of $\alpha$ and $\beta$ with $N=200$ and $\Lambda=10$
@@
$h$ $\mathrm{Err}_{h}(N)$ Order $\mathrm{Err}_{h}(N)$ Order $\mathrm{Err}_{h}(N)$ Order
@@
Figure 4: Numerical performance for the 1-D homogeneous problem with nonsmooth initial (4.2) at $t=0.4$ with $h=1/2^{7}$, $\alpha=0.6$, $\beta=0.75$, and $\Lambda=10$
Figure 5: Reference solution for the 1-D homogeneous problem with nonsmooth initial data (4.2) at different times $t$, with $h=1/2^{7}$, $\alpha=0.6$, $\beta=0.75$, and $\Lambda=1000$, $t_{0}=0.001$
Table 3: Temporal errors $\mathrm{Err}_{\tau}(N)$ at $t=0.4$ for the 2-D homogeneous problem with nonsmooth initial data (4.3), shown for different values of $\alpha$ and $\beta$ with $h=1/2^{7}$ and $\Lambda=10$
@@
Table 4: Temporal errors $\mathrm{Err}_{\tau}(N)$ at $t=0.1$ for the 2-D homogeneous problem with nonsmooth initial data (4.3), shown for different values of $\alpha$ and $\beta$ with $h=1/2^{7}$ and $\Lambda=10$
@@
Table 5: Spatial errors $\mathrm{Err}_{h}$ at $t=0.4$ for the 2-D homogeneous problem with nonsmooth initial data (4.3), computed for different values of $\alpha$ and $\beta$ with $N=200$ and $\Lambda=10$
@@
$h$ $\mathrm{Err}_{h}(N)$ Order $\mathrm{Err}_{h}(N)$ Order $\mathrm{Err}_{h}(N)$ Order
@@
(a) $t=0.001$
(b) $t=0.02$
(c) $t=0.04$
(d) $t=0.1$
Figure 6: Reference solutions for the 2-D homogeneous problem with nonsmooth initial value (4.3) at different times, with $\alpha=0.6$, $\beta=0.75$, and $\Lambda=1000$, $t_{0}=0.001$
Tables 1 and 3, along with Figure 5, demonstrate that the CIM achieves spectral convergence and high accuracy in the temporal direction. Furthermore, Tables 2 and 5 confirm that the spatial convergence order of the proposed algorithm is $\mathcal{O}(h^{2})$ for various values of $\alpha$ and $\beta$. These results are in full agreement with the theoretical analysis established in Theorem 3.
Moreover, compared with other algorithms, such as those in [22], our method attains superior precision. In particular, for the 2-D case, our algorithm delivers enhanced accuracy at a comparable spatial computational cost. Additionally, as evidenced by the comparison between Table 3 and Table 4, the computational expense of our algorithm remains independent of the final time $t$, since $t$ is treated as a parameter. This underscores a key advantage of our approach: it consistently maintains high accuracy even for problems with nonsmooth initial data.
@@ 4.3 Nonhomogeneous Problems with Vanishing Initial Values in 1-D and 2-D
Consider a special case of problem (1.1). In the 1-D case, the initial value is $u_{0}(x)=0$, with the source term
$f(x,t)=\frac{1}{\Gamma(2-\beta)}x(1-x)t^{1-\beta}+2t+\frac{2}{\Gamma(\alpha+2)}t^{\alpha+1},$ (4.4)
and the exact solution as
u(x,t)=tx(1x).u(x,t)=tx(1-x).
@@
Table 6: Temporal errors $\mathrm{Err}_{\tau}(N)$ at $t=0.6$ for the 1-D nonhomogeneous problem with source term (4.4), shown for different values of $\alpha$ and $\beta$ with $h=1/2^{9}$ and $\Lambda=10$
@@
20 5.4991e-04 4.1436e-04 3.0991e-04
40 4.1327e-08 2.7470e-08 1.6601e-11
60 (0.4,0.25) 2.4262e-12 (0.5,0.5) 1.4740e-12 (0.6,0.75) 1.2733e-12
80 1.1209e-12 2.4822e-13 1.8756e-12
100 8.9353e-13 6.9383e-13 2.4069e-13
Table 7: Spatial errors $\mathrm{Err}_{h}$ at $t=0.6$ for the 1-D nonhomogeneous problem with source term (4.4), computed for different values of $\alpha$ and $\beta$ with $N=200$ and $\Lambda=10$
@@
$h$ $\mathrm{Err}_{h}(N)$ Order $\mathrm{Err}_{h}(N)$ Order $\mathrm{Err}_{h}(N)$ Order
  $1/2^{5}$ 4.7089e-07 9.9536e-08 6.7171e-07
$2^{6} 1.1$773e-07 2.0000 2.4885e-08 1.9999 1.6793e-07 2.0000
$1/2^{7}$ 2.9453e-08 1.9990 6.1939e-09 2.0063 4.1962e-08 2.0007
$1/2^{8}$ 7.3008e-09 2.0123 1.5226e-09 2.0243 1.0551e-08 1.9918
(a) Absolute errors at $t=0.6$ with $\alpha=0.4$, $\beta=0.25$, and $\Lambda=10$
(b) Absolute errors at $t=0.6$ for different values of $\Lambda$, with $\alpha=0.4$ and $\beta=0.25$
Figure 7: Numerical performance for the 1-D nonhomogeneous problem with source term (4.4)
Table 8: Temporal errors $\mathrm{Err}_{\tau}(N)$ at $t=0.6$ for the 2-D nonhomogeneous problem with source term (4.5), shown for different values of $\alpha$ and $\beta$ with $h=1/2^{9}$ and $\Lambda=10$
@@
Table 9: Spatial errors $\mathrm{Err}_{h}$ at $t=0.6$ for the 2-D nonhomogeneous problem with source term (4.5), computed for different values of $\alpha$ and $\beta$ with $N=200$ and $\Lambda=10$
@@
$h$ $\mathrm{Err}_{h}(N)$ Order $\mathrm{Err}_{h}(N)$ Order $\mathrm{Err}_{h}(N)$ Order
@@ 4.4 Nonhomogeneous Problem with Nonsmooth Solution in the 1-D Case
$f(x,t)=1+\frac{\Gamma(\frac{7}{6})}{\Gamma(\frac{7}{6}-\beta)}x(1-x)t^{\frac{1}{6}-\beta}+2t^{\frac{1}{6}}+\frac{2\Gamma(\frac{7}{6})}{\Gamma(\alpha+\frac{7}{6})}t^{\alpha+\frac{1}{6}},$
@@
$\widehat{f}(x,z)=z^{-1}+\Gamma\left(\frac{7}{6}\right)\left(x(1-x)z^{\beta-\frac{7}{6}}+2z^{-\frac{7}{6}}+2z^{-\alpha-\frac{7}{6}}\right).$
@@
Table 10: Temporal errors $\mathrm{Err}_{\tau}(N)$ at $t=0.5$ for the 1-D nonhomogeneous problem with nonsmooth solution (4.6), shown for different values of $\alpha$ and $\beta$ with $h=1/2^{10}$ and $\Lambda=5$
@@
20 1.7717e-04 1.9252e-04 1.7580e-04
40 4.8009e-09 5.5776e-09 4.5635e-09
60 (0.25,0.4) 1.1870e-12 (0.5,0.6) 4.1117e-12 (0.75,0.8) 1.0706e-12
80 5.2034e-13 2.0441e-12 1.3599e-12
100 1.0798e-12 1.7649e-12 6.5225e-13
(a) Absolute errors at $t=0.5$ for different values of $\Lambda$, with $\alpha=0.25$ and $\beta=0.4$
(b) Absolute errors at different times $t$ for $\Lambda=15$, with $\alpha=0.75$ and $\beta=0.8$
Figure 8: Numerical performance for the 1-D nonhomogeneous problem with nonsmooth solution (4.6)
@@ 5 Conclusions
6 Acknowledgement
This work was partially supported by the National Natural Science Foundation of China under grants 42450192, 12071343, 11701416; Natural Science Basic Research Program of Shaanxi (Program No. 2025JC-YBMS-012), and basic research fund of Tianjin University under grant 2025XJ21-0010.

### v2

@@ 4.1 Scalar Problem
$\left\{\begin{aligned} &{{}_{0}^{C}}\!D_{t}^{\beta}u(t)+u(t)+\int_{0}^{t}\kappa_{\alpha}(t-s)u(s)\,ds=f(t),\quad t>0,\\ &u(0)=1,\end{aligned}\right$. (4.1)
@@
(a) The absolute errors for different $\Lambda$ at $t=0.5$ with $\alpha=0.2$, $\beta=0.77$.
(b) The absolute errors at different $t$ for $\Lambda=10$ with $\alpha=0.2$, $\beta=0.77$.
(c) The absolute errors for different $\alpha$ at $t=0.5$ with $\beta=0.77$, $\Lambda=10$.
(d) The absolute errors for different $\beta$ at $t=0.5$ with $\alpha=0.2$, $\Lambda=10$.
Figure 3: Numerical performances of the CIM for problem (4.1).
@@ 4.2 Homogeneous Problems with Nonsmooth Initial Values in 1-D and 2-D
Consider the homogeneous problem with nonsmooth initial values for problem (1.1). In the 1-D case, the initial value is given by
@@
with source term $f(x,t)\equiv 0$. In the 2-D case, the initial value is taken as
@@
and the source term is $f(x,y,t)\equiv 0$.
To evaluate the numerical performance, we use the numerical solution computed with $N=200$ and $h=1/2^{7}$ as the reference solution. The numerical results at $t=0.4$ are presented in Tables 1-5 and Figures 5. Additionally, Figures 5 and 6 illustrate the behaviors of reference solution at different times, which also confirm that the evolutionary behavior of the solution is influenced by the regularity of the initial condition, as discussed in Subsection 3.1. It also indirectly indicates that our algorithm is capable of accurately capturing this instantaneous evolutionary process near $t=0$. For a better observation of the solution’s variation, the parameters are taken as $t_{0}=0.001$ and $\Lambda=1000$ in these two figures.
Table 1: Temporal errors $\mathrm{Err}_{\tau}(N)$ at $t=0.4$ for the 1-D homogeneous problem with nonsmooth initial data (4.2), shown for different values of $\alpha$ and $\beta$ with $h=1/2^{7}$ and $\Lambda=10$.
@@
Table 2: Spatial errors $\mathrm{Err}_{h}(t)$ at $t=0.4$ for the 1-D homogeneous problem with nonsmooth initial data (4.2), computed for different values of $\alpha$ and $\beta$ with $N=200$ and $\Lambda=10$.
@@
$h$ $\mathrm{Err}_{h}(t)$ Order $\mathrm{Err}_{h}(t)$ Order $\mathrm{Err}_{h}(t)$ Order
@@
Figure 4: Numerical performance for the 1-D homogeneous problem with nonsmooth initial (4.2) at $t=0.4$ with $h=1/2^{7}$, $\alpha=0.6$, $\beta=0.75$, and $\Lambda=10$.
Figure 5: Reference solutions for the 1-D homogeneous problem with nonsmooth initial data (4.2) at different times, with $h=1/2^{7}$, $\alpha=0.6$, $\beta=0.75$, and $\Lambda=1000$, $t_{0}=0.001$.
Table 3: Temporal errors $\mathrm{Err}_{\tau}(N)$ at $t=0.4$ for the 2-D homogeneous problem with nonsmooth initial data (4.3), shown for different values of $\alpha$ and $\beta$ with $h=1/2^{7}$ and $\Lambda=10$.
@@
Table 4: Temporal errors $\mathrm{Err}_{\tau}(N)$ at $t=0.1$ for the 2-D homogeneous problem with nonsmooth initial data (4.3), shown for different values of $\alpha$ and $\beta$ with $h=1/2^{7}$ and $\Lambda=10$.
@@
Table 5: Spatial errors $\mathrm{Err}_{h}(t)$ at $t=0.4$ for the 2-D homogeneous problem with nonsmooth initial data (4.3), computed for different values of $\alpha$ and $\beta$ with $N=200$ and $\Lambda=10$.
@@
$h$ $\mathrm{Err}_{h}(t)$ Order $\mathrm{Err}_{h}(t)$ Order $\mathrm{Err}_{h}(t)$ Order
@@
(a) $t=0.001$.
(b) $t=0.02$.
(c) $t=0.04$.
(d) $t=0.1$.
Figure 6: Reference solutions for the 2-D homogeneous problem with nonsmooth initial value (4.3) at different times, with $\alpha=0.6$, $\beta=0.75$, and $\Lambda=1000$, $t_{0}=0.001$.
Tables 1 and 3, along with Figure 5, demonstrate that the CIM achieves spectral convergence and high accuracy in the temporal direction. Furthermore, Tables 2 and 5 confirm that the spatial convergence order of the proposed algorithm is $O(h^{2})$ for various values of $\alpha$ and $\beta$. These results are in full agreement with the theoretical analysis established in Theorem 3.
Moreover, compared with other algorithms, such as those in [22], our method attains superior precision. In particular, for the 2-D case, our algorithm delivers enhanced accuracy at a comparable spatial computational cost. Additionally, as evidenced by the comparison between Tables 3 and 4, the computational expense of our algorithm remains independent of the final time $t$, since $t$ is treated as a parameter. This underscores a key advantage of our approach: it consistently maintains high accuracy even for problems with nonsmooth initial data.
@@ 4.3 Nonhomogeneous Problems with Vanishing Initial Values in 1-D and 2-D
In the 1-D case, we take the initial value $u_{0}(x)\equiv 0$, with the source term
$f(x,t)=\frac{1}{\Gamma(2-\beta)}x^{3}(1-x)t^{1-\beta}+(12x^{2}-6x)\left(t+\frac{1}{\Gamma(\alpha+2)}t^{\alpha+1}\right),$ (4.4)
thus the exact solution is
$u(x,t)=tx^{3}(1-x).$
@@
Table 6: Temporal errors $\mathrm{Err}_{\tau}(N)$ at $t=0.6$ for the 1-D nonhomogeneous problem with source term (4.4), shown for different values of $\alpha$ and $\beta$ with $h=1/2^{10}$ and $\Lambda=10$.
@@
10 4.5381e-03 4.1416e-03 3.8911e-03
20 5.7582e-05 4.7460e-05 5.6301e-05
30 (0.4,0.25) 5.6110e-07 (0.5,0.5) 1.4720e-07 (0.6,0.75) 2.2743e-07
40 1.3154e-08 2.4810e-08 1.8756e-08
60 1.0070e-08 3.0381e-08 2.4069e-08
Table 7: Spatial errors $\mathrm{Err}_{h}(t)$ at $t=0.6$ for the 1-D nonhomogeneous problem with source term (4.4), computed for different values of $\alpha$ and $\beta$ with $N=80$ and $\Lambda=10$.
@@
$h$ $\mathrm{Err}_{h}(t)$ Order $\mathrm{Err}_{h}(t)$ Order $\mathrm{Err}_{h}(t)$ Order
  $1/2^{5}$ 6.4321e-06 7.9102e-06 6.1210e-06
$1/2^{6}$ 1.6542e-06 1.9610 2.0583e-06 1.9457 1.5938e-06 1.9472
$1/2^{7}$ 1.7421e-07 2.0015 2.2184e-07 2.0001 1.7240e-07 2.0003
$1/2^{8}$ 1.9823e-08 2.0450 2.5089e-08 2.0673 1.9543e-08 2.0581
(a) Absolute errors at $t=0.6$ with $\alpha=0.4$, $\beta=0.25$, and $\Lambda=10$.
(b) Absolute errors at $t=0.6$ for different values of $\Lambda$, with $\alpha=0.4$ and $\beta=0.25$.
Figure 7: Numerical performances for the 1-D nonhomogeneous problem with source term (4.4).
Table 8: Temporal errors $\mathrm{Err}_{\tau}(N)$ at $t=0.6$ for the 2-D nonhomogeneous problem with source term (4.5), shown for different values of $\alpha$ and $\beta$ with $h=1/2^{9}$ and $\Lambda=10$.
@@
Table 9: Spatial errors $\mathrm{Err}_{h}(t)$ at $t=0.6$ for the 2-D nonhomogeneous problem with source term (4.5), computed for different values of $\alpha$ and $\beta$ with $N=200$ and $\Lambda=10$.
@@
$h$ $\mathrm{Err}_{h}(t)$ Order $\mathrm{Err}_{h}(t)$ Order $\mathrm{Err}_{h}(t)$ Order
@@ 4.4 Nonhomogeneous Problem with Nonsmooth Solution in the 1-D Case
$f(x,t)=1+\frac{\Gamma(\frac{7}{6})}{\Gamma(\frac{7}{6}-\beta)}x^{3}(1-x)t^{\frac{1}{6}-\beta}+(12x^{2}-6x)\left(t^{\frac{1}{6}}+\frac{\Gamma(\frac{7}{6})}{\Gamma(\alpha+\frac{7}{6})}t^{\alpha+\frac{1}{6}}\right),$
@@
$\widehat{f}(x,z)=z^{-1}+\Gamma\Big(\frac{7}{6}\Big)\left(x^{3}(1-x)z^{\beta-\frac{7}{6}}+(12x^{2}-6x)\left(z^{-\frac{7}{6}}+z^{-\alpha-\frac{7}{6}}\right)\right).$
@@
Table 10: Temporal errors $\mathrm{Err}_{\tau}(N)$ at $t=0.5$ for the 1-D nonhomogeneous problem with nonsmooth solution (4.6), shown for different values of $\alpha$ and $\beta$ with $h=1/2^{10}$ and $\Lambda=5$.
@@
10 1.6714e-03 1.9751e-03 1.7548e-03
20 2.8019e-05 3.0776e-05 2.0855e-05
30 (0.25,0.4) 1.4872e-07 (0.5,0.6) 1.3044e-07 (0.75,0.8) 1.7584e-07
40 1.2034e-08 2.0471e-08 1.2758e-08
60 1.9078e-08 1.3649e-08 1.1572e-08
(a) Absolute errors at $t=0.5$ for different values of $\Lambda$, with $\alpha=0.75$ and $\beta=0.8$.
(b) Absolute errors at different times for $\Lambda=15$, with $\alpha=0.25$ and $\beta=0.4$.
Figure 8: Numerical performances for the 1-D nonhomogeneous problem with nonsmooth solution (4.6).